If there are two parallel plates that are separated by a distance of 0.01 m and they have a potential difference of 15000 V, find the electric field strength between the plates.
Answer #1

The relation between V and E is

V = - E d

E = - V/d

= -15000/0.01

= -1500*103 V/m

magnitude of electric field strength is 1500*103 V/m